The healthcare provider orders paclitaxel​ (Taxol) 135 mg/m2 to be administered IV over 24 hours. The​ client's body surface area is 1.62m2. The medication is supplied 6​ mg/mL. How many mL would you​ administer?

A.1,312.2 mL

B.36.5 mL

C.0.14 mL

D.13.9 mL
Answer #1

ANSWER

* Calculation of volume (ml) of drug required.

Formula

Volume (ml) = Desired dose/Dose in hand * Quantity

Here

Ordered dose = 135 mg/m2

BSA(m2) of the patient = 1.62 m2

So

Desired dose (D) = 135 * 1.62

= 218.7 mg

Dose in hand (H) = 6 mg

Quantity (Q) = 1 ml

As per above formula

Volume (ml) of drug required = D/H *Q

= 218.7/6 * 1

= 36.45 * 1

B = 36.5 ml